a) For each of the following t-accounts, calculate the missing amount and enter it in the space provided. Assume that the beginning and ending balances are on the normal side of the account; that is, the side on which the balance increases.

Service Fee Revenue
3/1 Beginning balance         6,000
3/6 Transaction A         3,000
3/28 Transaction C ? 3/22 Transaction B         3,500
Ending balance       10,500

b) For each of the following t-accounts, calculate the missing amount and enter it in the space provided. Assume that the beginning and ending balances are on the normal side of the account; that is, the side on which the balance increases.

Cash
1/1 Beginning balance ? 1/16 Transaction C         4,500
1/3 Transaction A         2,500 1/20 Transaction D         2,200
1/15 Transaction B         4,000 1/29 Transaction F         3,500
1/28 Transaction E         3,400
Ending balance         9,800

c) For each of the following t-accounts, calculate the missing amount and enter it in the space provided. Assume that the beginning and ending balances are on the normal side of the account; that is, the side on which the balance increases.

Accounts Payable
6/4 Transaction A         1,500 6/1 Beginning balance         3,000
6/8 Transaction C         2,000 6/6 Transaction B ?
6/29 Transaction D         3,400
Ending balance         6,700

d) For each of the following t-accounts, calculate the missing amount and enter it in the space provided.

Inventory
4/1 Beginning balance       40,000 4/9 Transaction A         8,000
4/15 Transaction B         5,000 4/28 Transaction C       25,000
Ending balance ?

e) For each of the following t-accounts, calculate the missing amount and enter it in the space provided.

Taxes Payable
2/1 Beginning balance       30,000
2/6 Transaction A       25,000
2/28 Transaction C ? 2/22 Transaction B         3,500
Ending balance       40,500
Answer #1

Part A

Transaction C = beginning balance + Transaction A + Transaction B – ending balance = 6000+3000+3500-10500 = $2000

Part B

Beginning balance = ending balance + Transaction C + Transaction D + Transaction F - Transaction A - Transaction B - Transaction E = 9800+4500+2200+3500-2500-4000-3400 = $10100

Part C

Transaction B = ending balance+ Transaction A + Transaction C - Beginning balance - Transaction D = 6700+1500+2000-3000-3400 = $3800

Part D

Ending balance = Beginning balance+Transaction B-Transaction A - Transaction C = 40000+5000-8000-25000 = $12000

Part E

Transaction C = beginning balance + Transaction A + Transaction B – ending balance = 30000+25000+3500-40500 = $18000